A pitcher throws a 0.2kg baseball. IF the ball starts from rest and leaves the pitcher's hand a velocity of 30m/s, how much work is done on the ball? Show complete solution.


1
Expert's answer
2021-12-20T10:31:36-0500

Given:

"m=0.2\:\rm kg"

"v_i=0\:\rm m/s"

"v_f=30\:\rm m/s"


The work done is equal to change in kinetic energy

"W=\frac{mv_f^2}{2}-\frac{mv_i^2}{2}=\rm\frac{30^2*0.2}{2}-\frac{0^2*0.2}{2}=90\: J"